# Condensing and Extracting Against Online Adversaries

### v2

@@ 1.4.3 Connection to oNOSF Sources
1.4.4 Previous Results Interpreted in oNOSF source context
Previous impossibility results can be interpreted in the context of extracting / condensing from uniform oNOSF sources . For instance, collective coin flipping impossibility results of [ben-or_collective_1989] imply extraction impossibility results for uniform ${\left(g,\ell,n\right)}$-oNOSF sources when $n=1$. They imply:
Corollary 1.8.
Similarly, we observe that the notion of collective sampling is equivalent to 00-error condensing. Hence, lower bounds of [ggl98] imply zero-error condensing lower bounds for uniform ${\left(g,\ell,n\right)}$-oNOSF sources when $n=1$. Formally:
Corollary 1.9.
1.4.5 $\varepsilon$-Collective Sampling
Since collective sampling lower bounds show that for any protocol, 00-error condensing beyond rate $g/\ell$ is impossible, one can naturally ask whether condensing with small error $\varepsilon$ is possible. We call this problem $\varepsilon$-collective sampling, where the goal is to output a distribution which is $\varepsilon$-close to a distribution where every output has small probability.
Interpreted this way, this is exactly what protocols arising out of our condensers for uniform oNOSF sources provide: Using 1, when each player has access to $10^{4}$ random bits, there exists a simple protocol that can handle $0.49\ell$ corrupt players such that the players can collectively sample a distribution over $m=O(\ell)$ bits which is $2^{-\Omega(\ell)}$-close to having entropy $0.99m$. As far as we are aware, such a protocol is not implied by any other previous protocol. Most previous protocols are obtained through leader election protocols, which do not seem useful here since the leader has access to only constant number of bits.
We similarly obtain explicit protocols using 3 for the case when each player has access to $n\geq 2^{\omega(\ell)}$ bits.
1.4.6 Collective Coin Flipping and Sampling with Weak Random Sources
A natural extension to collective coin flipping and sampling in the full information model is when all players only have access to weak source of randomness (that are independent from each other) instead of true uniform randomness. This question was first studied by [gsv05]. [klrz08network] used network extractor protocol to transform weak random sources of each player into independent private random sources. This way, after using the network extraction protocol, players can follow the usual collective coin flipping / sampling protocol. [gsz21network] improved the network extraction protocol using two-source non-malleable extractors.
Using our ${\left(g,\ell,n,k\right)}$-oNOSF source condensers, we obtain alternative, simple $\varepsilon$-collective sampling protocols in the setting where players have access to weak sources of randomness. We obtain such an existential protocol using 2, and explicit protocol using Corollary 1.3.
2 Proof Overview
Our proof overview begins by outlining our new existential results for condensers in Section 2.1 that is able to handle even constant block length. Next, we present our explicit condenser results in Section 2.2 before discussing of our low-entropy to uniform oNOSF source conversion in Section 2.3. We present the main ideas behind our results regarding online influence and extractor lower bounds in Section 2.4. In Section 2.5, we overview our extractor constructions for oNOBF and oNOSF sources, that is based on a general transformation from leader election protocols.
@@ 2.1 Existence of oNOSF Condensers for All $\ell$ and $n$
Here we sketch the proof of 1. This result states that when $g=0.51\ell$ and $n=1000$, there exists a condenser $\mathsf{Cond}$ for uniform ${\left(g,\ell,n\right)}$-oNOSF sources so that the output entropy rate is $0.99$, the number of output bits is $m=O(\ell+\log(1/\varepsilon))$, and the error of the condenser is $\varepsilon$ where $\varepsilon\leq 2^{-\Omega(\ell)}$ is arbitrary.
@@
Proof sketch for 1.
Let $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{\ell})$ be such a source. Let $\mathbf{Y}_{1}\sim(\{0,1\}^{n})^{0.5\ell}$ be the source obtained by concatenating the first $0.5\ell$ blocks of $\mathbf{X}$. Since $0.51\ell$ blocks are good, there exist at least $0.01\ell$ uniform blocks in $\mathbf{Y}_{1}$. We treat $\mathbf{Y}_{1}$ as a single distribution over $n\ell$ bits with min-entropy $\geq 0.01\ell n$.
Let $\mathbf{Y}_{2}\sim\{0,1\}^{0.5\ell}$ be the source obtained by concatenating $1$ bit from each of the last $0.5\ell$ blocks of $\mathbf{X}$. Once again, since $0.51\ell$ blocks are good, there exist at least $0.01\ell$ uniform bits in $\mathbf{Y}_{1}$.
We will use the following seeded condenser:
@@
Our condenser $\mathsf{Cond}$ will output $\mathsf{sCond}(\mathbf{Y}_{1},\mathbf{Y}_{2})$. Observe that not only is $\mathbf{Y}_{2}$ not uniform, there could be as many as $0.49\ell$ “bad bits” in $\mathbf{Y}_{2}$ that can depend on $\mathbf{Y}_{1}$. To remedy this, we use the well known fact that the behavior of such adversarial $\mathbf{Y}_{2}$ cannot be far worse than the behavior if $\mathbf{Y}_{2}$ were uniform. In particular, say if $\mathbf{Y}_{2}$ were uniform then the output entropy and error are $k$ and $\varepsilon$ respectively. Then for the actual $\mathbf{Y}_{2}$, the output entropy will be $k-0.49\ell$ and error will be $\varepsilon\cdot 2^{0.49\ell}$. See Lemma 4.9 for the formal statement.
@@
Hence, we finally obtain that our seeded condenser will output $0.01\ell n+O(\ell)$ bits and will have output entropy $m-\Delta$ where $\Delta=O(\ell)$. Hence, if $n$ is a large enough constant, our output entropy rate, $\frac{m-\Delta}{m}$, will be $\geq 0.99$ as desired.
@@ 2.2 Explicit Condensers for Uniform oNOSF Sources
Here we sketch the proof of 3: we construct explicit condensers for uniform ${\left(0.5\ell+1,\ell,n\right)}$-oNOSF sources where $\ell$ is a constant and $n$ is arbitrarily growing.
Our construction will be similar to that of the existential construction of condensers from [CGR_seedless_condensers]. We will use the online nature of these sources to make more observations that will allow us to obtain an explicit construction of such condensers using explicit seeded extractors (see Definition 3.4 for a definition) as our primitive:
Proof sketch for 3.
Let $\mathbf{X}$ be such a source. Let’s review the [CGR_seedless_condensers] construction at a high level: they take the first $\left\lceil\ell/2\right\rceil$ blocks and treat them as a single entity. From the remaining blocks, they take first few bits of each of the blocks with the number of such bits geometrically decreasing per block and concatenate them to obtain a second entity. They then pass these two sources to an “output-light” two-source extractor to obtain their final output.
@@
The construction of [CGR_seedless_condensers] had to introduce the notion of “output-lightness” to deal with the case of odd number of blocks. For instance, say $\ell=5$. Then, their first entity is obtained by concatenating the first $3$ blocks and second entity by taking careful number of bits from the remaining $2$ blocks. To handle scenarios where the first $3$ blocks were uniform and last $2$ were bad, the output-lightness property was imposed on two-source extractors, something which we do not know how to explicitly construct.
Just reducing to even cases is not enough since the construction of [CGR_seedless_condensers] required low-error two-source extractors with excellent parameters, and we do not know how to explicitly construct them. We bypass this requirement by further exploiting the fact that our adversary is online.
@@
2.3 Converting Low-Entropy oNOSF Sources to Uniform oNOSF Sources
They key part of our proof for condensing from low-entropy oNOSF sources is a transformation from low-entropy oNOSF sources to uniform oNOSF sources . Here, we sketch the proof for our transformation in Theorem 1.4 and compare it to that of [CGR_seedless_condensers]. Both these transformations rely on two-source extractors (see Definition 3.6 for definition) as a basic primitive.
Given a ${\left(g,\ell,n,k\right)}$-oNOSF source $\mathbf{X}=\mathbf{X}_{1},\dots,\mathbf{X}_{\ell}$, [CGR_seedless_condensers] uses excellent existential two-source extractors (such as from Lemma 6.6) to define output blocks $\mathbf{O}_{i}=2\mathsf{Ext}(\mathbf{X}_{1}\circ\cdots\circ\mathbf{X}_{i-1},\mathbf{X}_{i})$ for $i\in\{2,\dots,\ell\}$ and define their transformation as $f(\mathbf{X})=\mathbf{O}_{2},\dots,\mathbf{O}_{\ell}$. They show that $\mathbf{O}_{i}$ is a good block if: (1) $\mathbf{X}_{i}$ is a good block and (2) at least one block amongst $\mathbf{X}_{1},\dots,\mathbf{X}_{i-1}$ is a good block. They showed that such a good block will be uniform and independent of the blocks $\mathbf{O}_{2},\dots,\mathbf{O}_{i-1}$ and argued there will be $g-1$ such good output blocks. This indeed shows their output is a uniform ${\left(g-1,\ell-1,m\right)}$-oNOSF source. However, each of their output blocks has length $m=O\left(\frac{k}{\ell}\right)\leq O\left(\frac{n}{\ell}\right)$, and so they were not able to handle the case of $n=o(\ell)$. We improve on their construction by using a “sliding window” based technique to obtain a much better transformation that can even handle $n=\operatorname{poly}(\log(\ell))$.
Theorem 2.4 (Theorem 6.1 restated).
The parameter $d$ in our theorem statement above is the width of our sliding window. When we set $d=\ell$ we recover the analysis of [CGR_seedless_condensers]. The true advantage of our transformation emerges when $d$ is very small compared to $\ell$. For instance, when $g=0.51\ell,n=\operatorname{poly}(\log(\ell))$ and $k=\operatorname{poly}(\log(\ell))$, we set $d$ to be a large constant and conclude that the output distribution is a uniform ${\left(0.509\ell,\ell,\operatorname{poly}(\log(\ell)\right)}$-NOSF source.
Proof sketch of Theorem 2.4.
We first compute the number of good output blocks $g_{out}$. Let $j_{1},\dots,j_{g}$ be the indices of the good input blocks in $\mathbf{X}$ and $d_{i}=j_{i+1}-j_{i}$ be the gap between the $i$-th good block and the next ($i+1$)-th good block. If the gap $d_{i}$ is at most $d$, then $\mathbf{O}_{i+1}$ must be a good output block. So, $g_{out}$ is the number of $i$ such that $d_{i}\leq d$. Since $g\geq 0.51\ell$, such large gaps can’t appear too often and we can calculate that $g_{out}=g-\frac{\ell-g+2}{d}$ as desired.
Next, we show that the good output blocks are indeed uniform conditioned on all previous output blocks. With this, we will obtain that the output distribution will be uniform ${\left(g_{out},\ell-1,m\right)}$-oNOSF source as desired. Let $i$ be the index of a good output block. We want to show that $\mathbf{O}_{i}$ is uniform conditioned on $\mathbf{O}_{1},\dots,\mathbf{O}_{i-1}$. To do this, we first observe that any input block contributes to at most $d+1$ good output blocks. This means that$(\mathbf{X}_{i-d}\circ\cdots\circ\mathbf{X}_{i-1})$, which has min-entropy at least $k$, loses at most $d\cdot m$ min-entropy conditioned on fixing $\mathbf{O}_{1},\dots,\mathbf{O}_{i-1}$. Moreover, $\mathbf{X}_{i}$ still remains uniform and independent of $(\mathbf{X}_{i-d}\circ\cdots\circ\mathbf{X}_{i-1})$ when fixing these previous output blocks. Hence, the output of the two-source extractor will indeed be uniform as desired. ∎
2.4 Online Influence and Extractor Lower Bounds
In this subsection, we provide a brief overview of our results regarding online influence and sketch how they imply extractor lower bounds against oNOBF sources . We also contrast online with the established notion of influence for Boolean functions. For any function $f:\{0,1\}^{n}\to\{0,1\}$, define the function $e(f)(x)=(-1)^{f(x)}$.
A Poincaré inequality and extractor lower bounds
Theorem 2.5 (Theorem 7.5 restated).
The proof of Theorem 2.5 is based on techniques from the Fourier analysis of Boolean functions. [^6] We give a very brief recap of necessary notions from Fourier analysis of Boolean functions in Section 7.2. The following key result implies Theorem 2.5 in a straightforward way.
Lemma 2.6 (Lemma 7.7 restated).
The above bound is established using the following Fourier analytic characterization of online influence.
Claim 2.7 (7.8 restated).
The proof of the above result is mostly a standard Fourier analytic computation and we refer the reader to Section 7 for more details.
Influence vs Online Influence
We exactly exhibit such a separation via the non-monotone address function $\operatorname{Addr}_{\ell}:\{0,1\}^{\log\ell+\ell}\to\{0,1\}$ which considers its first $\log\ell$ bits as an index in $\{1,\dots,\ell\}$ and then outputs the value of the chosen index. It is easy to show (as we do in Lemma 7.12) that the first $\log\ell$ bits of $\operatorname{Addr}_{\ell}$ have no online influence, while the remaining bits have online influence of $O\left(\frac{1}{\ell}\right)$. This is in contrast to the well known result of [kahn_influence_1988] showing that, for a balanced function such as $\operatorname{Addr}_{\ell}$, there must exist a bit with influence at least $\Omega\left(\frac{\log\ell}{\ell}\right)$.
2.5 Extractors via Leader Election Protocols
Suppose $\pi$ is an $(r-1)$-round leader election protocol over $\ell$ players where in each round, each player sends $1$ bit and with the guarantee that if there are at most $\delta\ell$ bad players, then a good player is chosen as leader with probability $1-\epsilon$. Suppose $\mathbf{X}$ is an ${\left(g,\ell r\right)}$-oNOBF source, where $g\geq lr-\delta\ell$. We simply partition the bits of $\mathbf{X}$ into chunks $\mathbf{X}_{1},\mathbf{X}_{2},\ldots,\mathbf{X}_{r}$, where each $X_{i}$ is on $\ell$ bits, and simulate the protocol $\pi$ by using the $j$’th bit of $\mathbf{X}_{i}$ as the message of the $j$’th player in round $i$, for all $1\leq j\leq\ell$ and $11\leq i\leq r-$ 1. At the end of this simulation suppose $j^{*}\in[\ell]$ is the chosen leader. Then we output the $j^{*}$’th bit of $\mathbf{X}_{r}$ as the output of the extractor.
Briefly, the reason that the above is a valid simulation of $\pi$ is the fact that the value of any bad bit in this online setting just depends on bits that appear before it, which is allowed in the leader election protocol (where in round $i$, the message of a bad player can be any function of the messages in the same round or previous rounds). The correctness of the extractor now follows from the fact that since the number of bad players (i.e., bad bits in $\mathbf{X}$) is at most $\delta\ell$, it follows from the guarantee of the protocol that the chosen leader $j^{*}\in[\ell]$ is a good player with probability at least $1-\epsilon$, and in this case the $j^{*}$’th bit of $\mathbf{X}_{r}$ must be uniform.
We note here that in the usual definition of leader election protocols, the requirement is to select a good leader with constant probability, which is a weaker guarantee than what we need to instantiate the above plan. It turns out that we can combine leader election protocols from prior works, in particular from [feige_noncryptographic_1999] and [alonnaorcoin], to construct protocols with the stronger guarantee we require. We refer the reader to Section 9 for more details on the construction of our leader election protocols.
2.6 Organization
In the remainder of our paper, we give some preliminaries in Section 3 before moving on to our core results. Section 4 details our proofs for the existence of seedless condensers for oNOSF sources for all regimes of $\ell$ and $n$, while Section 5 provides proofs for our explicit constructions of condensers. Next, Section 6 shows how to handle converting low-entropy oNOSF sources to uniform oNOSF source for a broader range of parameters. In Section 7 we introduce the notion of online influence and use it to provide an extraction lower bound for oNOBF sources . In Section 8, we present our explicit constructions of extractors for oNOBF and oNOSF sources using a connection to leader election protocols. In Section 9, we explicitly construct the required leader election protocols. We discuss some open questions in Section 10.
In Appendix A, we consider a natural local variant of oNOSF sources and show that it is straightforward to extract from such sources using existing extractors for small-space sources.

### v3

@@ 2.3 Existence of oNOSF Condensers for All $\ell$ and $n$
Here we sketch the proof of Theorem 3. This result states that when $g=0.51\ell$ and $n=1000$, there exists a condenser $\mathsf{Cond}$ for uniform $\left(g,\ell,n\right)$}-oNOSF sources so that the output entropy rate is $0.99$, the number of output bits is $m=O(\ell+\log(1/\varepsilon))$, and the error of the condenser is $\varepsilon$ where $\varepsilon\leq 2^{-\Omega(\ell)}$ is arbitrary.
@@
Proof sketch for Theorem 3.
Let $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{\ell})$ be such a source. Let $\mathbf{Y}_{1}\sim(\{0,1\}^{n})^{0.5\ell}$ be the source obtained by concatenating the first $0.5\ell$ blocks of $\mathbf{X}$. Since $0.51\ell$ blocks are good, there exist at least $0.01\ell$ uniform blocks in $\mathbf{Y}_{1}$. We treat $\mathbf{Y}_{1}$ as a single distribution over $n\ell$ bits with min-entropy $\geq 0.01\ell n$. Let $\mathbf{Y}_{2}\sim\{0,1\}^{0.5\ell}$ be the source obtained by concatenating $1$ bit from each of the last $0.5\ell$ blocks of $\mathbf{X}$. Once again, since $0.51\ell$ blocks are good, there exist at least $0.01\ell$ uniform bits in $\mathbf{Y}_{2}$. We will use the following seeded condenser:
@@
Our condenser $\mathsf{Cond}$ will output $\mathsf{sCond}(\mathbf{Y}_{1},\mathbf{Y}_{2})$. Observe that not only is $\mathbf{Y}_{2}$ not uniform, there could be as many as $0.49\ell$ “bad bits” in $\mathbf{Y}_{2}$ that can depend on $\mathbf{Y}_{1}$. To remedy this, we use the well known fact that the behavior of such adversarial $\mathbf{Y}_{2}$ cannot be far worse than the behavior if $\mathbf{Y}_{2}$ were uniform. Concretely, suppose if $\mathbf{Y}_{2}$ were uniform and the output entropy and error were $k$ and $\varepsilon$. Then for the actual $\mathbf{Y}_{2}$, the output entropy will be $k-0.49\ell$ and error will be $\varepsilon\cdot 2^{0.49\ell}$. See Lemma 7.9 for the formal statement.
@@
Hence, we finally obtain that our seeded condenser will output $0.01\ell n+O(\ell)$ bits and will have output entropy $m-\Delta$ where $\Delta=O(\ell)$. Hence, if $n$ is a large enough constant, our output entropy rate, $\frac{m-\Delta}{m}$, will be $\geq 0.99$ as desired. ∎
@@
2.4 Online Influence and Extractor Lower Bounds
In this subsection, we provide a brief overview of our results regarding online influence and sketch how they imply extractor lower bounds against oNOBF sources . We also contrast this with the established notion of influence for Boolean functions. For any function $f:\{0,1\}^{n}\to\{0,1\}$, define the function $e(f)(x)=(-1)^{f(x)}$.
A Poincaré inequality and extractor lower bounds
Theorem 2.11 (Theorem 10.5 restated).
The proof of Theorem 2.11 is based on techniques from the Fourier analysis of Boolean functions. [^11] We give a very brief recap of necessary notions from Fourier analysis of Boolean functions in Section 10.2. The following key result implies Theorem 2.11. We refer the reader to Section 10 for more details.
Lemma 2.12 (Lemma 10.7 restated).
For any $f:\{0,1\}^{\ell}\to\{0,1\}$ and $i\in[{\ell}]$,
$\operatorname{\mathbf{oI}}_{i}(f)^{2}\leq\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(S)^{2}\leq\operatorname{\mathbf{oI}}_{i}(f).$
Influence vs Online Influence
We exactly exhibit such a separation via the address function $\operatorname{Addr}_{\ell}:\{0,1\}^{\log\ell+\ell}\to\{0,1\}$ which considers its first $\log\ell$ bits as an index in $\{1,\dots,\ell\}$ and then outputs the value of the chosen index. It is easy to show (as we do in Lemma 10.12) that the first $\log\ell$ bits of $\operatorname{Addr}_{\ell}$ have no online influence, while the remaining bits have online influence of $O\left(\frac{1}{\ell}\right)$. This is in contrast to the well known result of [KKL88] showing that, for a balanced function such as $\operatorname{Addr}_{\ell}$, there must exist a bit with influence at least $\Omega\left(\frac{\log\ell}{\ell}\right)$.
2.5 Extractors via Leader Election Protocols
Suppose $\pi$ is an (r1)(r-1)-round leader election protocol over $\ell$ players where in each round, each player sends $1$ bit and with the guarantee that if there are at most $\delta\ell$ bad players, then a good player is chosen as leader with probability $1-\epsilon$. Suppose $\mathbf{X}$ is an $\left(g,\ell r\right)$}-oNOBF source, where $g\geq lr-\delta\ell$. We simply partition the bits of $\mathbf{X}$ into chunks $\mathbf{X}_{1},\mathbf{X}_{2},\ldots,\mathbf{X}_{r}$, where each $X_{i}$ is on $\ell$ bits, and simulate the protocol $\pi$ by using the $j$’th bit of $\mathbf{X}_{i}$ as the message of the $j$’th player in round $i$, for all $1\leq j\leq\ell$ and $1\leq i\leq r-1$. At the end of this simulation suppose $j^{*}\in[\ell]$ is the chosen leader. Then we output the $j^{*}$’th bit of $\mathbf{X}_{r}$ as the output of the extractor.
Briefly, the reason that the above is a valid simulation of $\pi$ is the fact that the value of any bad bit in this online setting just depends on bits that appear before it, which is allowed in the leader election protocol (where in round $i$, the message of a bad player can be any function of the messages in the same round or previous rounds). The correctness of the extractor now follows from the fact that since the number of bad players (i.e., bad bits in $\mathbf{X}$) is at most $\delta\ell$, the guarantee of the protocol ensures that the chosen leader $j^{*}\in[\ell]$ is a good player with probability at least $1-\epsilon$, and in this case the $j^{*}$’th bit of $\mathbf{X}_{r}$ must be uniform.
We note here that in the usual definition of leader election protocols, the requirement is to select a good leader with constant probability, which is a weaker guarantee than what we need to instantiate the above plan. It turns out that we can combine leader election protocols from prior works, in particular from [Fei99] and [AN93], to construct protocols with the stronger guarantee we require. We refer the reader to Section 9 for more details on the construction of our leader election protocols.
2.6 Organization
We give some preliminaries in Section 4 before moving on to our core results. Section 5 provides proofs for our explicit condenser constructions with small block length, and Section 6 shows how to handle converting low-entropy oNOSF sources to uniform oNOSF source for a wider range of parameters. Next, Section 7 details our proof for the existence of seedless condensers for oNOSF sources for all regimes of $\ell$ and $n$. In Section 8, we present our explicit constructions of extractors for oNOBF and oNOSF sources using a connection to leader election protocols. Then, in Section 9, we explicitly construct the required leader election protocols. Finally, we introduce the notion of online influence in Section 10 and use it to provide an extraction lower bound for oNOBF sources . We discuss some open questions in Section 11.
In Appendix B, we consider a natural local variant of oNOSF sources and show that it is straightforward to extract from such sources using existing extractors for small-space sources.
@@ 2.7 Explicit Condensers for Uniform oNOSF Sources
Here we sketch the proof of Theorem 2: we construct explicit condensers for uniform $\left(0.5\ell+1,\ell,n\right)$}-oNOSF sources where $\ell$ is a constant and $n$ is arbitrarily growing.
Our construction will be similar to that of the existential construction of condensers from [CGR24]. We will use the online nature of these sources to make more observations that will allow us to obtain an explicit construction of such condensers using explicit seeded extractors (see Definition 4.5 for a definition) as our primitive:
Proof sketch for Theorem 2.
Let $\mathbf{X}$ be such a source. Let’s review the [CGR24] construction at a high level: they take the first $\left\lceil\ell/2\right\rceil$ blocks and treat them as a single entity. From the remaining blocks, they take first few bits of each of the blocks with the number of such bits geometrically decreasing per block and concatenate them to obtain a second entity. They then pass these two sources to an “output-light” two-source extractor to obtain their final output.
@@
The construction of [CGR24] had to introduce the notion of “output-lightness” to deal with the case of odd number of blocks. For instance, say $\ell=5$. Then, their first entity is obtained by concatenating the first $3$ blocks and second entity by taking careful number of bits from the remaining $2$ blocks. To handle scenarios where the first $3$ blocks were uniform and last $2$ were bad, the output-lightness property was imposed on two-source extractors, something which we do not know how to explicitly construct.
Just reducing to even cases is not enough since the construction of [CGR24] required low-error two-source extractors with excellent parameters, and we do not know how to explicitly construct them. We bypass this requirement by further exploiting the fact that our adversary is online.
@@
3 Application to Collective Coin Flipping and Collective Sampling
3.0.1 Background
Say there are $\ell$ players who have a common broadcast channel and want to jointly perform a task such as collectively flipping a coin. Some $b$ players out of them are “bad” and want to deter the task. We assume the bad players are computationally unbounded so cryptographic primitives are of no use. We further assume that each player has private access to uniform randomness. [BL89] initiated the study of this model and aptly termed this task as “collective coin flipping.”
The simplest way to collectively flip a coin would be for all the players to initially agree on a function $f:\{0,1\}^{\ell}\to\{0,1\}$, then synchronously broadcast one random bit $r_{i}$, and to finally agree on the output being $f(r_{1},\dots,r_{\ell})$. However, synchronizing broadcasts is hard, and it could be that the bad players set their output as function of the bits of the good players. [KKL88] showed that no function $f$ can handle more than $O\left(\frac{\ell}{\log\ell}\right)$ corruptions.
One way to allow for more corruptions (almost linear) among players is to consider “protocols” that allow more rounds of communication. In particular, a protocol can be thought of as a tree where each vertex represents a “round” where in every round the following happens: all good players sends their bits, then all bad players send their bits as a function of the bits of the good players, and they jointly compute a function of these bits. Depending on the outcome of the function, everyone branches on one branch in this tree. Furthermore, every leaf is labeled with final outcomes (say 0 or $1$) and, once a leaf is reached, that is the outcome that everybody agrees on. [GGL98] initiated the study of protocols where the outcomes are from a larger range and where the bad players are trying to minimize the largest probability of any outcome. They called this problem “collective sampling.” For a formal definition, see Section 4.4.
3.0.2 Known Results
[BL89] showed that for protocols with outcomes $\{0,1\}$, $b$ bad players can always ensure that some outcome occurs with probability at least $\frac{1}{2}+\frac{b}{2\ell}$. [AN93] first constructed a protocol that can handle a linear number of corruptions. Follow-up works tried to reduce the number of rounds in this protocol where, in some settings, players were allowed to send more than one bit per round [RZ01, Fei99].
[GGL98] showed that for all collective sampling protocols and all outcomes, there exists a way for $b$ bad players to coordinate and ensure that an outcome that happens without corruption with probability $p$, now happens with probability $p^{1-(b/n)}\geq p\left(1+\frac{b}{n}\log(1/p)\right)$. Nearly matching collective sampling protocols were constructed by [GGL98, SV08, GVZ06]. For an overview of further results and bounds, see [Dod06].
3.0.3 Connection to oNOSF Sources
The problem of extracting or condensing from oNOSF sources can be seen as special cases or variants of collective coin flipping and collective sampling that provide very simple protocols. For instance, suppose one has an extractor or condenser $f$ for $\textrm{uniform }{\left(g,\ell,n\right)}\textrm{-oNOSF sources}$. Then, consider a protocol where all $\ell$ players take turns and output $n$ random bits. The agreed final outcome is $f$ applied on these $\ell n$ bits. This leads to protocols that are structurally much simpler since players don’t have to carefully compute whose turn it is to go in various rounds and can obliviously prepare for their turn.
3.0.4 Previous Results Interpreted in oNOSF source context
Previous impossibility results can be interpreted in the context of extracting / condensing from uniform oNOSF sources . For instance, collective coin flipping impossibility results of [BL89] imply extraction impossibility results for uniform $\left(g,\ell,n\right)$}-oNOSF sources when $n=1$. They imply:
Corollary 3.1.
Similarly, we observe that the notion of collective sampling is equivalent to 0-error condensing. Hence, lower bounds of [GGL98] imply zero-error condensing lower bounds for uniform $\left(g,\ell,n\right)$}-oNOSF sources when $n=1$. Formally:
Corollary 3.2.
3.0.5 $\varepsilon$-Collective Sampling
Since collective sampling lower bounds show that for any protocol, 0-error condensing beyond rate $g/\ell$ is impossible, one can naturally ask whether condensing with small error $\varepsilon$ is possible. We call this problem $\varepsilon$-collective sampling, where the goal is to output a distribution which is $\varepsilon$-close to a distribution where every output has small probability.
Interpreted this way, this is exactly what protocols arising out of our condensers for uniform oNOSF sources provide: Using Theorem 3, when each player has access to $10^{4}$ random bits, there exists a simple protocol that can handle $0.49\ell$ corrupt players such that the players can collectively sample a distribution over $m=O(\ell)$ bits which is $2^{-\Omega(\ell)}$-close to having entropy $0.99m$. As far as we are aware, such a protocol is not implied by any other previous protocol. Most previous protocols are obtained through leader election protocols, which do not seem useful here since the leader has access to only constant number of bits.
We similarly obtain explicit protocols using Theorem 2 for the case when each player has access to $n\geq\operatorname{poly}(\log(\ell)/\varepsilon)$ many bits.
3.0.6 Collective Coin Flipping and Sampling with Weak Random Sources
A natural extension to collective coin flipping and sampling in the full information model is when all players only have access to weak source of randomness (that are independent from each other) instead of true uniform randomness. This question was first studied by [GSV05]. [KLRZ08] used network extractor protocol to transform weak random sources of each player into independent private random sources. This way, after using the network extraction protocol, players can follow the usual collective coin flipping / sampling protocol. [GSZ21] improved the network extraction protocol using two-source non-malleable extractors.
Using our $\left(g,\ell,n,k\right)$}-oNOSF source condensers, we obtain alternative, simple $\varepsilon$-collective sampling protocols in the setting where players have access to weak sources of randomness. We obtain such an existential protocol using Theorem 4, and explicit protocol using Corollary 1.3.